OpenStudy (anonymous):

the slope of the line is 3. find the point slope of the line. (6,8) is the answer is y - 6 = -3(x-8) ?

OpenStudy (anonymous):

The slope is 3, not -3. Also, you plugged in the point in the opposite direction, you put the x coordinate in the place where the y was supposed to go, and vice-versa.Point slope is y-y1 = m(x-x1). If you plug that all in, you get y-8 = 3(x-6).